If positive number is 5 less than another positive number. Six times lesser number minus 3 times the greater number is 3. Find the 2 positive numbers.
Answer by JulietG(1812)

We're first told that the two numbers are:
x [1st number] and (x+5)[2nd number]
The second sentence tells us this:
6(x) - (3)(x+5) = 3
or
6x - (3x+15) = 3
which becomes
6x - 3x - 15 = 3
Add 15 to each side:
6x - 3x = 3 + 15
Add/subtract:
3x = 18
Divide each side by 3:
x = 6
Let's plug it back in to make certain:
"If positive number is 5 less than another positive number. Six times lesser number minus 3 times the greater number is 3."
If x is 6, and it is 5 less than the other number, the other number is 11.
(6*6)-(3*11) = 3
36 - 33 does indeed equal 3.
Therefore the two numbers are 6 and 11.